OpenStudy (anonymous):

Help for medal! Simplify (x2 + 7x + 12)(x2 + 3x + 5). x4 + 26x2 + 60 x4 + 21x2 + 60 x4+ 3x3 + 5x2 + 10x + 17 x4 + 10x3 + 38x2 + 71x + 60

OpenStudy (igreen):

Multiply all the terms in the first parenthesis to all the terms in the 2nd parenthesis. x^2 * x^2 = ? x^2 * 3x = ? x^2 * 5 = ? 7x * x^2 = ? 7x * 3x = ? 7x * 5 = ? 12 * x^2 = ? 12 * 3x = ? 12 * 5 = ?

OpenStudy (anonymous):

45x^2+105x+60

OpenStudy (igreen):

@essence.m Please do not give answers, it's against OpenStudy's Code-of-Conduct. We are here to guide the asker to the answer, so we need his/her participation. Please read the \(\bf\color{blue}{\underline{{\href{http:///openstudy.com/code-of-conduct}{Code-of-Conduct}}}}\) to learn more.

OpenStudy (anonymous):

x^2 * x^2 = x^4 x^2 x 3x = 3x ^3 x^2 * 5 = 5x^2 7x * x^2 = 7x^3 7x * 3x = I think 21x but i'm not sure. 7x * 5 = is it 35x? 12 * x^2 = 12x^2 12 * 3x = 36x 12 * 5 = 60

OpenStudy (igreen):

Nice work, just one thing. 7x * 3x doesn't equal 21x Remember there are two 'x's, so we have: 21x^2

OpenStudy (igreen):

So the whole thing is: x^4 + 3x^3 + 5x^2 + 7x^3 + 21x^2 + 35x + 12x^2 + 36x + 60 Combine the like terms

OpenStudy (igreen):

Can you do that? @AMYCARTER

OpenStudy (igreen):

Like terms have the same variables and exponents. Here are the like terms: 3x^3 + 7x^3 5x^2 + 21x^2 + 12x^2 35x + 36x

OpenStudy (anonymous):

Sorry when the question shows me not viewing it I'm still here I don't know why it does that. I'll combine them real quick...

OpenStudy (igreen):

Yeah, one of OS's 23985093285702938572309857 glitches.

OpenStudy (anonymous):

3x^3 + 7x^3 = 10x^3 5x^2 + 21x^2 +12x^2 = 38x^2 35x + 36x = 71x^2 I think

OpenStudy (anonymous):

haha @iGreen

OpenStudy (igreen):

Nice work, but don't change the variables. 35x + 36x = 71x So our final answer is: x^4 + 10x^3 + 38x^2 + 71x + 60

OpenStudy (anonymous):

Okay I'll remember next time. Thank you for your help @iGreen I understand this type of problem better now! :)

OpenStudy (igreen):

No problem.